# Rethinking Model Prototyping through the MedMNIST+ Dataset Collection

### v2

@@ Abstract
The integration of deep learning based systems in clinical practice is often impeded by challenges rooted in limited and heterogeneous medical datasets. In addition, prioritization of marginal performance improvements on a few, narrowly scoped benchmarks over clinical applicability has slowed down meaningful algorithmic progress. This trend often results in excessive fine-tuning of existing methods to achieve state-of-the-art performance on selected datasets rather than fostering clinically relevant innovations. In response, this work presents a comprehensive benchmark for the MedMNIST+ database to diversify the evaluation landscape and conduct a thorough analysis of common convolutional neural networks (CNNs) and Transformer-based architectures, for medical image classification. Our evaluation encompasses various medical datasets, training methodologies, and input resolutions, aiming to reassess the strengths and limitations of widely used model variants. Our findings suggest that computationally efficient training schemes and modern foundation models hold promise in bridging the gap between expensive end-to-end training and more resource-refined approaches. Additionally, contrary to prevailing assumptions, we observe that higher resolutions may not consistently improve performance beyond a certain threshold, advocating for the use of lower resolutions, particularly in prototyping stages, to expedite processing. Notably, our analysis reaffirms the competitiveness of convolutional models compared to ViT-based architectures emphasizing the importance of comprehending the intrinsic capabilities of different model architectures. Moreover, we hope that our standardized evaluation framework will help enhance transparency, reproducibility, and comparability on the MedMNIST+ dataset collection as well as future research within the field. Code is available at github/sdoerrich97.
1 Introduction
@@
In recent years, significant strides in deep learning (DL) have reshaped various domains, from image classification to natural language processing (Wang et al., 2023). This progress was driven by the development of increasingly sophisticated models, exemplified by architectures like the Transformer (Vaswani et al., 2017) for text or Vision Transformer (ViT) (Dosovitskiy et al., 2021) for images. Moreover, advanced training methodologies, including self-supervised contrastive methods such as CLIP (Radford et al., 2021) for image and text pairs, and DINO (Caron et al., 2021; Oquab et al., 2024) for image pairs, have enabled the training of complex models without the need for exhaustive labeling efforts. Simultaneously, there is an accumulating interest in integrating machine learning techniques into medical imaging, where DL models are approaching performance comparable to medical experts on specific tasks (Liu et al., 2019) and software applications are beginning to receive clinical certifications (Sendak et al., 2020). Despite this progress and the exponential growth of DL-related publications across various medical fields in the past few years (Kocak et al., 2023), the adoption of DL algorithms in daily clinical practice has been comparatively slow (Stacke et al., 2021).
@@
One major obstacle is the scarcity of appropriate datasets, often characterized by limited sample sizes and heterogeneous image acquisition conditions (Lafarge et al., 2017; Oksuz et al., 2020; Khan et al., 2022), thereby posing challenges to the generalizability of supervised DL algorithms. Ongoing progress in domain adaptation (DA) and domain generalization (DG) techniques aims to increase algorithmic robustness through aligning feature distributions (Li et al., 2021) or acquiring domain-invariant features (Li et al., 2018). However, the generalizability of these methods across diverse domains remains a significant challenge, constraining their real-world applicability Eche et al. (2021).
@@
In addition, there is a concerning trend in DL research towards prioritizing the adaptation and scaling of existing methodologies in order to achieve incremental performance improvements on influential benchmarks (Raji et al., 2021) rather than addressing clinically relevant needs (Varoquaux and Cheplygina, 2022). This trend is particularly pronounced in academic research, where the incentive structure often prioritizes quantity over relevance, leading to the incorporation of additional complexity into existing methodologies often at the expense of increased computational requirements (Janiesch et al., 2021). While benchmark datasets play a crucial role in coordinating machine learning research and facilitating standardized evaluations (Koch et al., 2021), overreliance on a handful of influential yet narrowly scoped benchmarks may stifle innovation and exacerbate inherent biases within these datasets such as the underrepresentation of certain demographic groups (Crawford and Paglen, 2021; Birhane and Prabhu, 2021). The latter, in particular, limits the applicability of current DL techniques across diverse patient populations, thereby impeding their real-world deployment (Norori et al., 2021). Instead, research endeavors should focus more on proposing new benchmarks to diversify the landscape, mitigate bias-induced challenges, and cover a broader range of real-world tasks. Rather than solely determining a winner based on state-of-the-art performance, benchmarking should promote understanding to drive impactful algorithmic development and alternative evaluation methods (Raji et al., 2021).
@@
Furthermore, the limitations of scaling alone are becoming increasingly evident, as larger models start to falter in model trustworthiness (Rae et al., 2021; Thoppilan et al., 2022) or performance on well-specified tasks (McKenzie et al., 2022). Nonetheless, there is a paramount trend of increasing hardware and compute requirements estimated via the total number of FLOPs, and the number of trainable parameters in deep learning architectures (Sevilla et al., 2022). This further impedes the application of these approaches in the clinical environment. Therefore, it is imperative to explore qualitative enhancements alongside quantitative scaling in DL research as called for by Goyal and Bengio (2022), particularly in the context of real-world medical applications.
@@
Research endeavors should prioritize the creation of larger and more diverse datasets and benchmarks, with a focus on incorporating additional inductive biases and fostering the continuous development of more sophisticated approaches. The recent emergence of foundation models exemplifies this direction. These models, pre-trained on extensive datasets, offer the potential to enhance performance by capturing intricate patterns and serving as a foundational basis for further fine-tuning (Bommasani et al., 2021). Existing works, building on top of these models, can be readily evaluated across diverse benchmarks due to their high transferability to new datasets and tasks, as well as their remarkable zero- or few-shot performance (Kirillov et al., 2023; Girdhar et al., 2023). This facilitates a more comprehensive assessment of these methods without necessitating extensive retraining.
@@
In this work, we aim to contribute to this effort by reassessing traditional DL models and training schemes, and presenting a new benchmark in the context of medical image classification. Our objective is to reevaluate commonly held assumptions regarding these methodologies, thereby enhancing and confirming comprehension of their inherent strengths and limitations as well as diversifying the benchmark landscape. Consequently, we will offer recommendations and insights for prototyping, model development, and deployment. To this end, we extend upon the existing MedMNIST v2 classification benchmark (Yang et al., 2023), using the recently introduced MedMNIST+ database  [^1] MedMNIST+ including official dataset splits: https://zenodo.org/records/10519652. MedMNIST v2 offers a collection of 12 distinct biomedical 2D datasets, ranging from Chest-X-ray to Dermatology, in a MNIST-like (Deng, 2012) resolution of $28\times 28$ pixels for medical image analysis. Its limitation to $28\times 28$ images represented a critical constraint for comprehensive method evaluation. However, this has been addressed with the introduction of MedMNIST+, extending the previous dataset collection with resolutions: $64\times 64$, $128\times 128$, and $224\times 224$ pixels. By systematically benchmarking a diverse array of baseline models and training paradigms, including selective convolutional and Transformer-based models using both end-to-end training and linear probing on this distinct multi-dimensional database, our goal is to provide critical insights into the strengths and weaknesses of these techniques. Furthermore, we investigate the integration of the k-nearest neighbors ($k$-NN) classifier into the feature space of these models, aiming to enhance computational efficiency and interpretability. Our primary aim is to re-investigate whether compute-intensive architectures are always necessary, whether higher resolution input consistently improves model performance, and whether end-to-end training is always optimal. Additionally, we want to foster greater transparency, reproducibility, and comparability in future research endeavors within the domain of medical image analysis. Key contributions of our work include:
@@
2 Method
2.1 Model Selection
Our selection of model architectures encompasses a diverse array of both convolutional and Transformer-based networks. Among the chosen convolutional models are well-established architectures such as VGG16 (Simonyan and Zisserman, 2014), AlexNet (Krizhevsky et al., 2012), ResNet-18 (He et al., 2015), DenseNet-121 (Huang et al., 2016), and EfficientNet-B4 (Tan and Le, 2019), all of which were pretrained on the ImageNet1k dataset (Russakovsky et al., 2014). In the domain of Transformers, we include the Vision Transformer (ViT) (Dosovitskiy et al., 2021), renowned for its exceptional performance across various image classification tasks, pretrained on ImageNet1k as well as the CLIP (Radford et al., 2021) and DINO (Caron et al., 2021) pretrained ViT variants. Acknowledging recent advancements in this area, our selection extends to adaptations of ViT such as EVA-02 (Fang et al., 2023) and the Segment Anything Model (SAM) (Kirillov et al., 2023). EVA-02 represents a series of efficiently optimized plain ViTs with moderate model sizes, employing bidirectional visual representations learned from a robust CLIP encoder. Conversely, SAM, initially conceived as a foundational model for image segmentation, is intentionally designed and trained to be promptable, thus facilitating zero-shot transferability to new image distributions and tasks, including image classification. For all ViT architectures, we opted for the base backbone with a patch size of 16 (i.e. ViT-B/16). With the exception of the AlexNet model obtained from the torchvision library, all models were sourced from the "Pytorch Image Models (timm)" library (Wightman, 2019) at Huggingface.222timm at Hugginface: https://huggingface.co/timm [^3] Model identifier convolution: vgg16.tv_in1k, x resnet18.a1_in1k, x densenet121.ra_in1k,
xxii
efficientnet_b4.ra2_in1k
 [^4] Model identifier transformer: vit_base_patch16_224.augreg2_in21k_ft_in1k,
xxii
vit_base_patch16_clip_224.laion2b_ft_in12k_in1k,
xxii
eva02_base_patch16_clip_224.merged2b_s8b_b131k, x vit_base_patch16_224.dino,
xxii
samvit_base_patch16.sa1b
Further details regarding the employed backbone architectures, including parameter counts, the number of activations, Giga Multiply-Add Operations per Second (GMACs), and feature dimension before the last classification layer, are outlined in Table 1.
Table 1: Details of evaluated model architectures including parameter count, number of activations, Giga Multiply-Add Operations per Second (GMACs) and feature dimension before the last classification layer. The number of parameters and activations is given in million (M). Except for CLIP ViT-B/16 and EVA-02 ViT-B/16 (both CLIP pretrained), DINO ViT-B/16 (DINO pretrained) and SAM ViT-B/16 (pretrained on SA-1B), all models were pretrained on ImageNet1K.
Model Params (M) Activations (M) GMACs # Output Dimension
VGG1633footnotemark: 3 138.4 13.6 15.5 4096
AlexNet33footnotemark: 3 62.3 0.6 0.36 4096
ResNet-1833footnotemark: 3 11.7 2.5 1.8 512
DenseNet-12133footnotemark: 3 8 6.9 2.9 1024
EfficientNet-B433footnotemark: 3 19.3 34.8 3.1 1792
ViT-B/1644footnotemark: 4 86.6 16.5 16.9 768
CLIP ViT-B/1644footnotemark: 4 86.6 16.5 16.9 768
EVA-02 ViT-B/1644footnotemark: 4 86.3 16.5 16.9 768
DINO ViT-B/1644footnotemark: 4 85.8 16.5 16.9 768
SAM ViT-B/1644footnotemark: 4 89.7 1343.3 486.4 256
2.2 Training Pipeline
We adopt diverse training paradigms, encompassing both end-to-end training (i.e. training the whole model), and linear probing, where we solely train the classification head, while keeping the encoder frozen. Additionally, we explore the integration of the k-nearest neighbors ($k$-NN) classifier into the feature space of pre-trained models. Following Nakata et al. (2022) and Doerrich et al. (2024), the pre-trained image encoder initially extracts feature embeddings from the training set, which are then stored in an external database along with their corresponding labels. During inference, the image encoder generates a feature embedding for a given query image. Subsequently, the top-$k$ feature embeddings having highest similarity scores (i.e. lowest cosine distance) with the query embedding are retrieved from the training set along with their associated labels. The classification of the query image is afterward determined through a majority vote on these labels. This approach facilitates efficient classification without necessitating retraining of the classification head or the entire encoder, thereby enhancing computational efficiency, interpretability, and generalizability, with reduced dependence on hyperparameters. Given the substantial computational cost associated with training deep neural networks, particularly foundation models, the adoption of $k$-NN methods presents an efficient alternative to traditional training schemes.
The training regimen consisted of 100 epochs with early stopping based on the validation set. We employed the AdamW optimizer (Loshchilov and Hutter, 2017) with a learning rate of $0.0001$, along with a cosine annealing learning rate scheduler (Loshchilov and Hutter, 2016) with a single cycle. Each model was trained with a batch size of $64$, allowing for training on a single NVIDIA RTX™ A5000 GPU. For evaluations utilizing the k-nearest neighbors ($k$-NN) approach, we set $k$ to $11$, in line with Zhu et al. (2021), who demonstrated the suitability of $k>10$ for detecting noisy labels. To maintain compatibility with the pretrained models while preserving the inherent properties of individual resolutions, all image resolutions were padded to $224\times 224$ pixels using zero padding. This choice was motivated by Hashemi (2019), which demonstrated that zero-padding has no discernible effect on classification accuracy while significantly reducing training time compared to image resizing. With zero-padding, neighboring zero input units (pixels) do not activate their corresponding convolutional unit in the subsequent layer, resulting in decreased requirements for updating synaptic weights on outgoing links and ensuring robust feature preservation during image reshaping.
2.3 Loss Criterion and Evaluation Metrics
In line with the methodology outlined by Yang et al. (2023), we select the choice of loss criteria to suit the specific classification tasks associated with each dataset. For binary (BC) and multi-class classification (MC), as well as ordinal regression (OR) tasks, we utilize the Cross-Entropy (CE) loss function applied to the logits:
Furthermore, for the purpose of simplicity and standardization, consistent with Yang et al. (2023), we employ the same evaluation metrics including accuracy (ACC) and the area under the receiver operating characteristic curve (AUC) to assess the model’s ability to differentiate between classes.
3 Experiments and Results
3.1 Datasets
@@
The dataset selection employed in this work originates from MedMNIST v2 (Yang et al., 2023), initially introduced at a resolution of $28\times 28$ pixels and recently expanded by MedMNIST+ to four distinct image resolutions, namely $28\times 28$, $64\times 64$, $128\times 128$, and $224\times 224$. The collection comprises twelve 2D datasets that are curated from carefully selected sources, encompassing primary data modalities such as X-ray, OCT, ultrasound, CT, and electron microscope. Furthermore, these datasets cater to diverse classification tasks, including binary/multi-class, ordinal regression, and multi-label classification, spanning a wide range of dataset scales, ranging from 780 samples for Breast up to 236,386 for Tissue. The details of each dataset including data source, data modality, type of the classification task (plus number of classes) as well as the publicly available data splits, provided by MedMNIST11footnotemark: 1and forming a one-to-one correspondence with our benchmark, are described in Table 2. In addition, Figure 1 illustrates the datasets side-by-side for all four image resolutions.
@@
Table 2: Dataset details including data source, data modality, type of the classification task (and number of classes) as well as data splits. (ML: Multi-Label, MC: Multi-Class, BC: Binary-Class, OR: Ordinary Regression).
Dataset Source Modality Task # Train / Val / Test
(# Classes)
Blood Acevedo et al. (2020) Blood Cell Microscope MC (8) $11,959$ / $1,712$ / $3,421$
Breast Al-Dhabyani et al. (2020) Breast Ultrasound BC (2) $546$ / $78$ / $156$
Chest Wang et al. (2017) Chest X-Ray ML-BC (2) 78,468 / 11,219 / 22,433
Derma Tschandl et al. (2018) Dermatoscope MC (7) 7,007 / 1,003 / 2,005
Codella et al. (2019)
OCT Kermany et al. (2018) Retinal OCT MC (4) 97,477 / 10,832 / 1,000
OrganA Bilic et al. (2023) Abdominal CT MC (11) 34,561 / 6,491 / 17,778
Xu et al. (2019)
OrganC Bilic et al. (2023) Abdominal CT MC (11) 12,975 / 2,392 / 8,216
Xu et al. (2019)
OrganS Bilic et al. (2023) Abdominal CT MC (11) 13,932 / 2,452 / 8,827
Xu et al. (2019)
Path Kather et al. (2019) Colon Pathology MC (11) 89,996 / 10,004 / 7,180
Pneumonia Kermany et al. (2018) Chest X-Ray BC (2) 4,708 / 524 / 624
Retina Liu et al. (2022) Fundus Camera OR (5) 1,080 / 120 / 400
Tissue Ljosa et al. (2012) Kidney Cortex Microscope MC (8) 165,466 / 23,640 / 47,280
3.2 Benchmark Evaluation
@@
We report the performance assessment of each model selected from our diverse pool, following the methodology outlined in Section 2.2. Our evaluation encompasses all datasets, image resolutions, and training schemes described earlier. The summary of average accuracy (ACC) and area under the receiver operating characteristic curve (AUC) across all datasets is provided in Table 3. In addition, detailed performance evaluations for each dataset individually can be found in the appendix, spanning Tables 6 to 17. To ensure the reliability of our assessments for all datasets, we report the mean and standard deviation of ACC and AUC for three random seeds.
@@
As anticipated, end-to-end training yields the highest overall performance for all training schemes, and higher resolutions appear to enable all models across all training schemes to exhibit performance enhancements compared to lower resolutions. However, these enhancements begin to plateau when transitioning from inputs of $128\times 128$ pixels to $224\times 224$ pixels. Despite the increased data information, characterized by a fourfold increase in pixel count, all models and training schemes across all datasets show only marginal improvements or, in some cases, even worse overall performance. This correspondence is visually depicted in Figure 2, where the accuracy distributions of each model across all datasets are displayed for each training scheme and input resolution, respectively. Furthermore, the performance relationships of the models to each other for a specific training scheme and input resolution remain largely consistent. This observation challenges the common assumption that evaluations solely on higher image resolutions (e.g. above $200\times 200$ pixels) are deemed valid, while evaluations on lower input resolutions are generally considered less meaningful, since the performance trends appear to be resolution independent. This in turn supports the utilization of lower resolution inputs, particularly during the prototyping phase of model development, as they generally allow for faster processing speeds while demanding fewer computational resources.
@@
Table 3: Benchmark outcomes summarizing the average mean and standard deviation of accuracy (ACC) and area under the receiver operating characteristic curve (AUC) across all datasets for all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.
@@
3.3 Input Resolution Impact
We further investigate the effect of input resolution on model performance. Our objective is to determine how often model performance enhances with incremental increases in input resolution. Intuitively, higher input resolutions are expected to enable models to capture more intricate features, potentially leading to improved overall performance. However, as demonstrated in Section 3.2, this trend reaches a plateau around an input resolution of $128\times 128$ pixels for our underlying setting.
To validate this observation, we analyze the instances where a model’s performance surpasses that of the previous, lower resolution for each training scheme individually. Specifically, we compare the mean accuracy values across three different random seeds for the same model and training scheme between two resolutions. Figure 3 depicts this analysis for transitions from $28\times 28$ to $64\times 64$, $64\times 64$ to $128\times 128$, and $128\times 128$ to $224\times 224$ resolutions. We calculate the frequency of performance improvements per increase in input resolution across all 12 datasets, resulting in a maximum improvement count of 12 for each transition.
Our findings substantiate the observations from Section 3.2 to some extent. Overall, higher input resolutions lead to performance improvements across all models and training schemes. However, this improvement diminishes notably when transitioning from a $128\times 128$ to $224\times 224$ resolution, with superior performance observed only for a limited number of dataset instances. This trend is particularly pronounced for end-to-end trained convolutional models, whereas ViT-based models exhibit less sensitivity to input resolution variations. This discrepancy could be attributed to the specific design of the ViT architecture, which is tailored for $224\times 224$ pixel images, unlike convolutional models. Additionally, we observe that linear probing benefits the most from higher resolution images, with slight differences noted for the $k$-NN approach, potentially due to the pretraining with images of the same size. These results underscore that while input resolution impacts model performance, the effect is less significant than initially anticipated, with slight variations depending on the architecture used. This supports the utility of lower input resolutions at least during the prototyping phase.

### v3

@@ Abstract
The integration of deep learning based systems in clinical practice is often impeded by challenges rooted in limited and heterogeneous medical datasets. In addition, the field has increasingly prioritized marginal performance gains on a few, narrowly scoped benchmarks over clinical applicability, slowing down meaningful algorithmic progress. This trend often results in excessive fine-tuning of existing methods on selected datasets rather than fostering clinically relevant innovations. In response, this work introduces a comprehensive benchmark for the MedMNIST+ dataset collection, designed to diversify the evaluation landscape across several imaging modalities, anatomical regions, classification tasks and sample sizes. We systematically reassess commonly used Convolutional Neural Networks (CNNs) and Vision Transformer (ViT) architectures across distinct medical datasets, training methodologies, and input resolutions to validate and refine existing assumptions about model effectiveness and development. Our findings suggest that computationally efficient training schemes and modern foundation models offer viable alternatives to costly end-to-end training. Additionally, we observe that higher image resolutions do not consistently improve performance beyond a certain threshold. This highlights the potential benefits of using lower resolutions, particularly in prototyping stages, to reduce computational demands without sacrificing accuracy. Notably, our analysis reaffirms the competitiveness of CNNs compared to ViTs, emphasizing the importance of comprehending the intrinsic capabilities of different architectures. Finally, by establishing a standardized evaluation framework, we aim to enhance transparency, reproducibility, and comparability within the MedMNIST+ dataset collection. Code is available at github.com/sdoerrich97/rethinking-model-prototyping-MedMNISTPlus.
keywords:
benchmarking, prototyping recommendations, medical image classification, standardized evaluation framework, foundation models
Introduction
@@
In recent years, significant strides in deep learning (DL) have reshaped various domains, from image classification to natural language processing [1]. This progress was driven by the development of increasingly sophisticated models, exemplified by architectures like the Transformer [2] for text or Vision Transformer (ViT) [3] for images. Moreover, advanced training methodologies, including self-supervised contrastive methods such as CLIP [4] for image and text pairs, and DINO [5, 6] for image pairs, have enabled the training of complex models without the need for exhaustive labeling efforts. Simultaneously, there is an accumulating interest in integrating machine learning techniques into medical imaging, where DL models are approaching performance comparable to medical experts on specific tasks [7] and software applications are beginning to receive clinical certifications [8]. Despite this progress and the exponential growth of DL-related publications across various medical fields in the past few years [9], the adoption of DL algorithms in daily clinical practice has been comparatively slow [10].
@@
One major obstacle is the scarcity of appropriate datasets, often characterized by limited sample sizes and heterogeneous image acquisition conditions [11, 12, 13], thereby posing challenges to the generalizability of supervised DL algorithms. Ongoing progress in domain adaptation (DA) and domain generalization (DG) techniques aims to increase algorithmic robustness through aligning feature distributions [14] or acquiring domain-invariant features [15]. However, the generalizability of these methods across diverse domains remains a significant challenge, constraining their real-world applicability [16].
@@
In addition, there is a concerning trend in DL research towards prioritizing the adaptation and scaling of existing methodologies in order to achieve incremental performance improvements on influential benchmarks [17] rather than addressing clinically relevant needs [18]. This trend is particularly pronounced in academic research, where the incentive structure often prioritizes quantity over relevance, leading to the incorporation of additional complexity into existing methodologies often at the expense of increased computational requirements [19]. While benchmark datasets play a crucial role in coordinating machine learning research and facilitating standardized evaluations [20], overreliance on a handful of influential yet narrowly scoped benchmarks may stifle innovation and exacerbate inherent biases within these datasets such as the underrepresentation of certain demographic groups [21, 22]. The latter, in particular, limits the applicability of current DL techniques across diverse patient populations, thereby impeding their real-world deployment [23]. Instead, research endeavors should focus more on proposing new benchmarks to diversify the landscape, mitigate bias-induced challenges, and cover a broader range of real-world tasks. Rather than solely determining a winner based on state-of-the-art performance, benchmarking should promote understanding to drive impactful algorithmic development and alternative evaluation methods [17].
@@
Furthermore, the limitations of scaling alone are becoming increasingly evident, as larger models start to falter in model trustworthiness [24, 25] or performance on well-specified tasks [26]. Nonetheless, there is a paramount trend of increasing hardware and compute requirements estimated via the total number of FLOPs, and the number of trainable parameters in deep learning architectures [27]. This further impedes the application of these approaches in the clinical environment. Therefore, it is imperative to explore qualitative enhancements alongside quantitative scaling in DL research as called for by A. Goyal and Y. Bengio [28], particularly in the context of real-world medical applications.
@@
Research endeavors should prioritize the creation of larger and more diverse datasets and benchmarks, with a focus on incorporating additional inductive biases and fostering the continuous development of more sophisticated approaches. The recent emergence of foundation models exemplifies this direction. These models, pre-trained on extensive datasets, offer the potential to enhance performance by capturing intricate patterns and serving as a foundational basis for further fine-tuning [29]. Existing works, building on top of these models, can be readily evaluated across diverse benchmarks due to their high transferability to new datasets and tasks, as well as their remarkable zero- or few-shot performance [30, 31]. This facilitates a more comprehensive assessment of these methods without necessitating extensive retraining.
@@
In this work, we aim to contribute to this effort by reassessing traditional DL models and training schemes, and presenting a new benchmark in the context of medical image classification. Our objective is to reevaluate commonly held assumptions regarding these methodologies, thereby enhancing and confirming comprehension of their inherent strengths and limitations as well as diversifying the benchmark landscape. Consequently, we will offer recommendations and insights for prototyping, model development, and deployment. To this end, we extend upon the existing MedMNIST v2 classification benchmark [32], using the recently introduced MedMNIST+ database [33]. MedMNIST v2 offers a collection of 12 distinct biomedical 2D datasets, ranging from Chest-X-ray to Dermatology, in a MNIST-like [34] resolution of $28\times 28$ pixels for medical image analysis. Its limitation to $28\times 28$ images represented a critical constraint for comprehensive method evaluation. However, this has been addressed with the introduction of MedMNIST+, extending the previous dataset collection with resolutions: $64\times 64$, $128\times 128$, and $224\times 224$ pixels. By systematically benchmarking a diverse array of baseline models and training paradigms, including selective convolutional and Transformer-based models using both end-to-end training and linear probing on this distinct multi-dimensional database, our goal is to provide critical insights into the strengths and weaknesses of these techniques. Furthermore, we investigate the integration of the k-nearest neighbors ($k$-NN) classifier into the feature space of these models, aiming to enhance computational efficiency and interpretability. Given that most clinically validated and regulated systems currently rely on supervised deep learning, in particular CNN-based architectures [35, 36], we prioritize these widely adopted models in our analysis. Our primary aim is to re-investigate whether compute-intensive architectures are always necessary, whether higher resolution input consistently improves model performance, and whether end-to-end training is always optimal. Additionally, we want to foster greater transparency, reproducibility, and comparability in future research endeavors within the domain of medical image analysis. Key contributions of our work include:
@@
Experiments and Results
Datasets
@@
The dataset selection employed in this work originates from MedMNIST v2 [32], initially introduced at a resolution of $28\times 28$ pixels and recently expanded by MedMNIST+ [33] to four distinct image resolutions, namely $28\times 28$, $64\times 64$, $128\times 128$, and $224\times 224$. The collection comprises twelve 2D datasets that are curated from carefully selected sources, encompassing primary data modalities such as X-ray, OCT, ultrasound, CT, and electron microscope. Furthermore, these datasets cater to diverse classification tasks, including binary/multi-class, ordinal regression, and multi-label classification, spanning a wide range of dataset scales, ranging from 780 samples for Breast up to 236,386 for Tissue. The details of each dataset including data source, imaging modality, type of classification task (along with the number of classes), and the publicly available data splits, provided by MedMNIST [33] and forming a one-to-one correspondence with our benchmark, are described in Table 1. Additionally, we report Shannon’s Equitability for each split to quantify class imbalance. A visual comparison of all datasets across the four evaluated image resolutions is presented in Figure 1.
@@
Table 1: Dataset details including data source, imaging modality, type of classification task (with number of classes), predefined data splits, and class imbalance measured using Shannon’s Equitability (ranging from 0 for total imbalance to 1 for perfect balance) for each split. (ML: Multi-Label, MC: Multi-Class, BC: Binary-Class, OR: Ordinary Regression).
Dataset Source Imaging Modality Task Number of Samples Class Imbalance
(# Classes) Train / Val / Test Train / Val / Test
Blood A. Acevedo et al. [37] Blood Cell Microscope MC (8) $11,959$ / $1,712$ / $3,421$ 0.96 / 0.96 / 0.96
Breast W. Al-Dhabyani et al. [38] Breast Ultrasound BC (2) $546$ / $78$ / $156$ 0.84 / 0.84 / 0.84
Chest X. Wang et al. [39] Chest X-Ray ML-BC (2) 78,468 / 11,219 / 22,433 0.26 / 0.26 / 0.26
Derma P. Tschandl et al. [40] Dermatoscope MC (7) 7,007 / 1,003 / 2,005 0.58 / 0.58 / 0.58
N. Codella et al. [41]
OCT D. S. Kermany et al. [42] Retinal OCT MC (4) 97,477 / 10,832 / 1,000 0.84 / 0.84 / 1.00
OrganA P. Bilic et al. [43] Abdominal CT MC (11) 34,561 / 6,491 / 17,778 0.96 / 0.95 / 0.96
X. Xu et al. [44]
OrganC P. Bilic et al. [43] Abdominal CT MC (11) 12,975 / 2,392 / 8,216 0.95 / 0.95 / 0.95
X. Xu et al. [44]
OrganS P. Bilic et al. [43] Abdominal CT MC (11) 13,932 / 2,452 / 8,827 0.93 / 0.96 / 0.94
X. Xu et al. [44]
Path J. N. Kather et al. [45] Colon Pathology MC (11) 89,996 / 10,004 / 7,180 0.99 / 0.99 / 0.96
Pneumonia D. S. Kermany et al. [42] Chest X-Ray BC (2) 4,708 / 524 / 624 0.82 / 0.82 / 0.95
Retina R. Liu et al. [46] Fundus Camera OR (5) 1,080 / 120 / 400 0.87 / 0.86 / 0.87
Tissue V. Ljosa et al. [47] Kidney Cortex Microscope MC (8) 165,466 / 23,640 / 47,280 0.87 / 0.87 / 0.87
Benchmark Evaluation
@@
We report the performance assessment of each model selected from our diverse pool, following the methodology outlined in Section Training Pipeline. Our evaluation encompasses all datasets, image resolutions, and training schemes described earlier. The summary of average accuracy (ACC) and area under the receiver operating characteristic curve (AUC) across all datasets is provided in Table 2. In addition, detailed performance evaluations for each dataset individually can be found in the appendix, spanning Supplementary Tables S1 to S12. To ensure the reliability of our assessments for all datasets, we report the mean and standard deviation of ACC and AUC for three random seeds.
@@
As anticipated, end-to-end training yields the highest overall performance for all training schemes, and higher resolutions appear to enable all models across all training schemes to exhibit performance enhancements compared to lower resolutions. However, these enhancements begin to saturate when transitioning from inputs of $128\times 128$ pixels to $224\times 224$ pixels. Despite the increased data information, characterized by a fourfold increase in pixel count, all models and training schemes across all datasets show only marginal improvements or, in some cases, even worse overall performance. This correspondence is visually depicted in Figure 2, where the accuracy distributions of each model across all datasets are displayed for each training scheme and input resolution, respectively. Furthermore, the performance relationships of the models to each other for a specific training scheme and input resolution remain largely consistent. This observation challenges the common assumption that evaluations solely on higher image resolutions (e.g. above $200\times 200$ pixels) are deemed valid, while evaluations on lower input resolutions are generally considered less meaningful, since the performance trends appear to be resolution independent. This in turn supports the utilization of lower resolution inputs, particularly during the prototyping phase of model development, as they generally allow for faster processing speeds while demanding fewer computational resources.
@@
Table 2: Benchmark outcomes summarizing the average mean and standard deviation of accuracy (ACC), for a fixed operating point of $0.5$, and area under the receiver operating characteristic curve (AUC) across all datasets for all training scheme-model-image resolution combinations, derived from three independent random seeds. Notably, the $k$-NN algorithm, devoid of a training phase, remains unaffected by the stochasticity inherent in model training, thus reporting only the total ACC value without standard deviation. Moreover, owing to its direct utilization of embeddings and labels for classification, $k$-NN does not furnish a reliable AUC score. The overall best result across all training schemes, models, and resolutions is highlighted with a background color; the best result per resolution across all training schemes and models is highlighted with underline; and the best result per training scheme and resolution is highlighted in bold.
@@
Input Resolution Impact
We further investigate the effect of input resolution on model performance. Our objective is to determine how often model performance enhances with incremental increases in input resolution. Intuitively, higher input resolutions are expected to enable models to capture more intricate features, potentially leading to improved overall performance. However, as demonstrated in Section Benchmark Evaluation, this trend reaches a saturation point around an input resolution of $128\times 128$ pixels for our underlying setting. To validate this observation, we analyze the instances where a model’s performance surpasses that of the previous, lower resolution for each training scheme individually. Specifically, we compare the mean accuracy values across three different random seeds for the same model and training scheme between two resolutions. Figure 3 depicts this analysis for transitions from $28\times 28$ to $64\times 64$, $64\times 64$ to $128\times 128$, and $128\times 128$ to $224\times 224$ resolutions. We calculate the frequency of performance improvements per increase in input resolution across all 12 datasets, resulting in a maximum improvement count of 12 for each transition.
Our findings substantiate the observations from Section Benchmark Evaluation to some extent. Overall, higher input resolutions lead to performance improvements across all models and training schemes. However, this improvement diminishes notably when transitioning from a $128\times 128$ to $224\times 224$ resolution, with superior performance observed only for a limited number of dataset instances. This trend is particularly pronounced for end-to-end trained convolutional models, whereas ViT-based models demonstrate less sensitivity to input resolution variations. This discrepancy could be attributed to the specific design of the ViT architecture, which is tailored for $224\times 224$ pixel images, unlike convolutional models. Additionally, we observe that linear probing benefits the most from higher resolution images, with slight differences noted for the $k$-NN approach, potentially due to the pretraining with images of the same size. These results underscore that while input resolution impacts model performance, the effect is less significant than initially anticipated, with slight variations depending on the architecture used. This supports the utility of lower input resolutions at least during the prototyping phase.